OpenStudy (anonymous):

alpha and beta are quadrant 1 angles with tan (alpha) = 8/15 and sec (beta) = 41/40. find sin (alpha + beta)

OpenStudy (anonymous):

Pythagorean theorem for any right triangle... and sin (alpha + beta) = sin alpha cos beta + sin beta cos alpha

OpenStudy (anonymous):

for alpha, your tan is 8/15, meaning opposite side o=8 and adjacent side a=15, therefore \[h=\sqrt{8^2+15^2}=\sqrt{64+225}=\sqrt{289}=17\]

OpenStudy (anonymous):

you can now solve sin alpha and cos alpha based on function definitions... sin alpha = opposite/hypotenuse cos alpha = adjacent/hypotenuse

OpenStudy (anonymous):

same procedure for the beta.... sec beta = hypotenuse/adjacent = 41/40.... which means opposite is missing, still you can solve it by Pythagorean Theorem... after solving opposite... you can solve for sin beta sin beta = opposite/hypotenuse for cos beta = 1/sec beta = 40/41

OpenStudy (anonymous):

once you get all of sin & cos of both alpha & beta...

OpenStudy (anonymous):

so I add 17 and 57.3

OpenStudy (anonymous):

no we'll use the identity \[\sin (\alpha+\beta)=\sin \alpha \cos \beta + \sin \beta \cos \alpha\]

OpenStudy (anonymous):

\[\sin \alpha = \frac{ 8 }{ 17 }\]

OpenStudy (anonymous):

ok...got that

OpenStudy (anonymous):

\[\cos \alpha = \frac{ 15 }{ 17 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\cos \beta = \frac{ 40 }{ 41 }\]

OpenStudy (anonymous):

41/40?

OpenStudy (anonymous):

for opposite side to angle beta.... we know h=41 and a=40 so \[o =\sqrt{41^2-40^2}=\sqrt{1681-1600}=\sqrt{81}=9\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

since sec beta = 1/cos beta...

OpenStudy (anonymous):

\[\sin \beta =\frac{ 9 }{ 41 }\]

OpenStudy (anonymous):

\[\sin(\alpha+\beta)=\frac{ 8 }{ 17 }\frac{ 40 }{ 41 }+\frac{ 9 }{ 41 }\frac{ 15 }{ 17 }\]

OpenStudy (anonymous):

you can proceed for simplification... :)

OpenStudy (anonymous):

do I multiply the first two numbers and then add them?

OpenStudy (anonymous):

Multiply first then add...

OpenStudy (anonymous):

great! you're the best.

OpenStudy (anonymous):

not really, just a Math enthusiast... :)